OpenStudy (lukecrayonz):

Differentiate f(x)=(8+x^3)^3-(4+x^6)^5

OpenStudy (anonymous):

power rule plus chain rule \[\frac{d}{dx}(8+x^3)^3=3(8+x^2)^2\times 3x^2\] second part is similar